OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned?

OpenStudy (anonymous):

i think you are supposed to solve \[x+y=100\]\[2x-3y=20\] for \(x\) and \(y\) the first equation is the same as \(y=100-x\) so the second equation becomes \[2x-3(100-x)=20\] can you solve this for \(x\)?

terenzreignz (terenzreignz):

If only to elaborate, satellite73 let x = number of men and y = number of women... and quoting the problem: "100 men and women" x + y = 100 "twice the number of men less three times the number of women is equal to 20" 2x - 3y = 20 And you have your two linear equations sufficient to solve for the value of x :)